# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v2

@@ 4.1 Verification with global power spectra
Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times. Even though this penalizes the ERA5-trained GC, Figs. 2c-d indicate that GC long-lead inferences are superior to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.
@@ 4.2 Verification against radiosondes
At ECCC, model developers often rely on forecast verification against ECMWF-HRES analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al. (2008), Privé et al. (2021)). Conversely, since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems. To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network. Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa are therefore used as reference for verification.
Two flavors of verification against radiosonde observations are performed based on the results shown in the previous section: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC. The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns (Mass et al., 2002).
The filter applied here is a spectral filter based on spherical harmonics. It approximately removes scales smaller than 1000 km (total wavenumber 40) and fully preserves scales larger than 2000 km (wavenumber 20). These cutoff wavenumbers well reflect the effective resolution of GC at forecast days 3 to 5 (See Figs. 2a and b). Note also that the same filter configuration is applied to GC and GDPS-CTL (more details on this filter can be found in Appendix A).
Figure 3: Verification for 120-hr forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter. The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression. Red and blue markers denote statistically significant differences in favor of GC and GDPS-CTL, respectively. Significance is computed using the $t$-test for bias and the $F$-test for STDE. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. A low pass filter (Appendix A) is applied on both GC and GDPS-CTL to obtain the low-resolution fields (e–h).
@@ 4.2.1 At full resolution
Comparing the forecasts at their native resolution against radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL. Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-hr forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig. 3a–d). The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature, where clear improvements can be seen with GC. In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression. Above 200 hPa, however, GC suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
To summarize the scores at different lead times, variables and vertical levels a single forecast quality index based on the fractional root-mean-square error (RMSE) change is defined as follows:
$f_{RMSE}=\left[\frac{RMSE(x_{1})}{RMSE(x_{2})}-1\right]\times 100\%,$ (3)
with $x_{1}$ and $x_{2}$ here representing GDPS-CTL and GC, respectively. Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite. The forecast quality index $f_{RMSE}$ was computed every 24 hr for each selected variable and level. The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 10 and 12.
@@
Results for boreal winter (see Fig. 4a) show that $f_{RMSE}$ is quite constant with lead times for each variable and layer. The largest improvements from GC are obtained in the low-level layer for all variables and decrease with altitude, turning into deteriorations in the high-level layer for geopotential heights and temperatures. GraphCast geopotential heights are also degraded with respect to GDPS-CTL in the mid-level layer at short lead times. Results for boreal summer are roughly similar (not shown). The pressure weighting approach adopted for the training of GC (see Fig. 6 in Lam et al. (2022)) is likely responsible for this vertical variation in performances. The differences in the upper-level layer between ECMWF-ERA5 and GDPS analyses used for GC initial conditions (e.g., different bias correction approaches for radiances, different data assimilated, distinct biases in the IFS and GEM model, etc.) could also explain some of the degradations observed in the upper-level layer when GC is initialized with the GDPS analysis. However, GC initialized with ECMWF-ERA5 has also been shown to increase the RMSE with respect to ECMWF-HRES above the tropopause when compared to analyses (see, e.g., Fig. 19 in Lam et al. (2022)).
@@
The region-wise variations in relative forecast performances can be examined by computing, as before, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits. In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs. 5a and c). In the high-level layer, GDPS-CTL performs better during each hemisphere’s respective winter. However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe. The largest improvements from GC in the mid- and high-level layers are found over the Tropics, which is consistent with the results of Lam et al. (2022) (see their Fig. 19).
@@
Figure 4: Changes in the forecast quality index (Eq. 3) against global radiosonde observations, averaged over 3 layers and 60 cases during boreal winter 2022, corresponding to forecasts at (a) high resolution and (b) low resolution. Upward-pointing red triangles indicate a reduction of the RMSE in GC forecasts with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. The size of the triangles varies linearly up to a value of 5.0%. When the index is greater than this threshold, the size is kept constant, but the rounded values are shown. Triangles are color-filled if the significances computed using an $F$-test exceeds the 90th percentile. The values on the left-hand side show the averaged indices over all the lead times and all the variables for each layer.
@@
Figure 5: Heatmap of the changes in the forecast quality index (Eq. 3) against global radiosonde observations for various geographical domains and atmospheric layers for (a, b) 60 cases of boreal winter 2022 and (c, d) 62 cases of boreal summer 2022. Verifications are presented for forecasts (a, c) at their full resolution and (b, d) after applying a low pass filter (Appendix A) on both GC and GDPS-CTL. Red shadings indicate a reduction of the RMSE by GC with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@ 4.2.2 At low resolution
At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs. 3e–h and 4b). This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 hr (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 hr (Fig. 5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig. 4a). This suggests that the improvements with GC for the first 72 hr at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
Degradations in GC, seen at full resolution over Northern Hemisphere during boreal winter (mostly coming from Europe) and Southern hemisphere during boreal summer, are enhanced at low resolution (compare Figs 5a,c, and 5b,d). There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig. 5d). The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig. 5), which confirms that the lower effective resolution of GC provides a significant advantage over the GDPS-CTL. Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers. This suggests that the GC forecasts contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
@@ 5.2 Implementation in GEM
To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales. For global simulations, such a filter should ideally employ a spherical harmonics-based spectral decomposition. However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost. As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids. Additionally, nudging is separately applied to the Yin and Yang grid solutions after every dynamics sub-step. It is important to note that, for the range of cut-off scales considered in this study, the response of the DCT-based filter (applied separately to the Yin-Yang grids) is found to be closely equivalent to the response of a spherical-harmonics-based filter applied to a global Gaussian field (see Appendix B).
Large scales that are retained with spectral nudging are defined through configurable parameters. Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter. With this method, the spectral filter relies on two different cut-off scale parameters denoted by $\lambda_{LS}$ and $\lambda_{SS}$ such that scales larger than $\lambda_{LS}$ are fully retained while those smaller than $\lambda_{SS}$ are completely eliminated. A partial filtering, similar to the one utilized by Husain et al. (2014), is applied for the scales in between $\lambda_{LS}$ and $\lambda_{SS}$. Further information on the DCT-based filter and its response is provided in Appendix B.
@@ 5.3 Optimal configuration
A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS. Results presented in section 6 are based on this optimal configuration which is referred to as GDPS-SN. Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
@@ 5.3.3 Nudging length scales
It was shown earlier in section 4 that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs. 1 and 2). The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging. However, improvements with such an approach over stationary cut-off scales are found to be negligible. Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km. This implies that scales larger than 2750 km are entirely retained, while those smaller than 2250 km are completely eliminated. The scales in between are subjected to partial filtering (see Fig. 19). This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs. 1 and 2).
@@ 6.1 Verification with global power spectra
The power spectra of the GDPS-SN forecast variables is indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing. The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq. 1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN (Fig. 7). This provides a solid proof-of-concept for the spectral nudging approach. The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs. 7a, b). This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4). Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL. This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
@@
Comparison of the spectral coherence (Eq. 2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs. 7c, d). These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d). The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration. Nudging the smaller scales would therefore bring the coherence on par with GC, particularly for short lead times, but it would result in variance deficiency, as discussed earlier.
Figure 7: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL. Additionally, for convenience of comparison, GC has been added in Fig. 7c and d.
@@
An illustrative example of the value added by the new hybrid system over GC is provided in Fig. 8. For temperature, which is explicitly nudged in GDPS-SN, the large-scale features of its transient-eddy anomalies tend to very closely resemble GC (Fig. 8a–c). Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are also found to be more aligned with GC than with GDPS-CTL (Fig. 8d–f). This is however not surprising as it reinforces the findings of previous studies on spectral nudging (Husain et al., 2014). Regarding fine scales, Fig. 8 clearly shows that GDPS-SN has similar level of variability as GDPS-CTL, which is not the case with GC as it clearly suffers from severe smoothing.
@@
Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature and (d–f) specific humidity, valid 22 February 2022 at 12z for GDPS-CTL (left), GC (middle) and GDPS-SN (right).

### v3

@@ 3.1 Verification with global power spectra
Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times. Even though this penalizes the ERA5-trained GraphCast, Figs. 2c-d indicate that GraphCast long-lead inferences are more skillful compared to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.
@@ 3.2 Verification against radiosondes
At ECCC, model developers often rely on forecast verification against ECMWF analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al. (2008), Privé et al. (2021)). Conversely, GraphCast’s training likely introduces some dependency on the ECMWF analysis system. To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to mainly rely on observations from the global radiosondes network for zonal wind ($U$), wind modulus ($|\mathbf{V_{H}}|$), geopotential height ($Z$), temperature ($T$), and dewpoint depression ($T–T_{d}$; where $T_{d}$ denotes dewpoint temperature) at mandatory pressure levels between 1000 and 100 hPa.
Two flavors of verification against radiosonde observations are performed based on the results shown in section 3a: (1) verification at full resolution, with the GDPS-CTL and GraphCast forecasts at their native resolutions, and (2) verification at a low resolution –- corresponding to the effective resolution of GraphCast (around day 3) – wherein both the GDPS-CTL and GraphCast forecasts are filtered to remove the variability at scales not adequately resolved by GraphCast. The latter is required because GraphCast’s smoothing artificially reduces the “double penalty” for misplaced patterns in the nominal resolution evaluation (Mass et al., 2002). Low resolution fields are computed using a spectral filter based on spherical harmonics. The response function of the filter reflects the effective resolution of GraphCast at forecast days 3 to 5 (Figs. 2a and b; details in appendix B).
Figure 3: Verification for 120-hr forecasts from GraphCast (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter. The error standard deviation (STDE; solid lines) and bias (mean forecast-minus-observation; dashed lines) are shown for (a,e) wind modulus, (b,f) geopotential heights, (c,g) temperature, and (d,h) dewpoint depression. Red and blue markers denote statistically significant differences in favor of GraphCast and GDPS-CTL, respectively. Significance is computed using the $t$-test for bias and the $F$-test for STDE. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. A low pass filter (appendix B) is applied on both GraphCast and GDPS-CTL to obtain the low-resolution fields (e–h).
@@ 3.2.1 At full resolution
Comparing the forecasts at their native resolutions against radiosonde observations reveals that GraphCast forecasts are generally closer to the observations than GDPS-CTL in the troposphere. Figures 3a–d show the corresponding results at day 5. While GraphCast shows clear improvements in temperature biases (dashed lines), it suffers from some noticeable deceleration within the troposphere, leading to deteriorations in wind modulus biases. In terms of error standard deviation (STDE, solid lines), however, GraphCast clearly shows large improvements over GDPS-CTL for wind modulus and dewpoint depression, although suffers from some deterioration in the mid-to-upper troposphere geopotential height. Above 200 hPa, GraphCast suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
To summarize the scores at different lead times, variables, and vertical levels, a single forecast quality index based on the change in root-mean-square error (RMSE) is defined as
$f_{RMSE}=\left[\frac{RMSE(x_{1})}{RMSE(x_{2})}-1\right]\times 100\%$ (3)
where $x_{1}$ and $x_{2}$ represent GDPS-CTL and GraphCast, respectively. Positive values of $f_{RMSE}$ indicate improved forecasts from GraphCast, while negative values imply degradations. A $f_{RMSE}$ was computed every 24 hr for each selected variable and level, with the results vertically averaged over three layers: low level (from 1000 to 850 hPa), mid level (from 700 to 300 hPa), and upper level (from 250 hPa to 100 hPa).
Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 11 and 13.
@@
Results for boreal winter show that $f_{RMSE}$ is quite constant with lead times for most variables and layers (Fig. 4a). The largest improvements from GraphCast are obtained in the low-level layer for all variables and decrease with altitude, turning into deteriorations in the upper-level layer for geopotential heights and temperatures. GraphCast geopotential heights are also degraded with respect to GDPS-CTL in the mid-level layer at short lead times, which is consistent with the error profile presented in Fig. 3b. Results for boreal summer are roughly similar (not shown). The pressure weighting approach adopted for the training of GraphCast (see Fig. 6 in Lam et al. 2022) is likely responsible for this vertical variation in performances. Systematic differences in the weakly constrained upper-level layer between ECMWF (used for fine-tuning) and GDPS analyses (used for initializing GraphCast) – e.g., different bias correction approaches for radiances, different data assimilated, distinct biases in the IFS and GEM model, etc. – could also explain some of the degradations observed in the upper-level layer. However, GraphCast initialized with ECMWF-ERA5 has also been shown to increase stratospheric RMSE with respect to ECMWF analyses (see, e.g., Fig. 19 in Lam et al. 2022).
@@
Regional variations in relative forecast performance can be examined by computing $f_{RMSE}$ on the six subdomains listed in Table 1. In low- and mid-level layers, GraphCast forecasts show smaller RMSE in every subdomain in both seasons (Figs. 5a and b). At upper levels, GDPS-CTL performs better during each hemisphere’s respective winter and over Asia during boreal summer. However, GraphCast’s poor performance during winter over the Northern Hemisphere is primarily originating from Europe and Asia. The largest improvements from GraphCast in the mid- and upper-level layers are found over the Tropics, as reported by Lam et al. (2022) (see their Fig. 19). This consistency in GraphCast’s regional performance with respect to its initialization with different analyses (ECMWF and GDPS analyses) is reassuring.
@@
Figure 4: Changes in the forecast quality index (Eq. 3) against global radiosonde observations, averaged over 3 layers and 60 cases during boreal winter 2022, corresponding to forecasts at (a) full resolution and (b) low resolution. Upward-pointing red triangles indicate a reduction of RMSE by GraphCast with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. The size of the triangles varies linearly up to a value of 5.0%. When the index is greater than this threshold, the size is kept constant, but the rounded values are shown. Triangles are color-filled if the significances computed using an $F$-test exceeds the 90th percentile. The values on the left-hand side show the averaged indices over all the lead times and all the variables for each layer.
@@
Figure 5: Heatmap of the changes in the forecast quality index (Eq. 3) against global radiosonde observations for various geographical domains and atmospheric layers averaged over all lead times for (a, c) 60 cases of boreal winter 2022, and (b, d) 62 cases of boreal summer 2022. Verifications are presented for forecasts (a, b) at their full resolution, and (c, d) after applying a low pass filter (appendix B) on both GraphCast and GDPS-CTL. Red shadings indicate a reduction of the RMSE by GraphCast with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@ 3.2.2 At low resolution
At low resolution, GraphCast forecast improvements over GDPS-CTL are reduced for all variables, levels, lead times, seasons, and geographical domains (Figs. 3e–h, 4b, and 5c,d). The change is particularly striking in the mid-level layer over the first 72 hr (compare Figs. 4a and b), with considerably smaller GraphCast forecast improvements than in the full-resolution case. The forecast improvements in low- and full-resolution context, however, become more aligned around 120 hr (Fig. 4b). This stands in contrast with a relatively steady improvement at all lead times found at full resolution (Fig. 4a), implying that these apparent short-range GraphCast forecast improvements, with respect to GDPS-CTL, originate primarily from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty. At low resolution, as expected, both GraphCast and GDPS-CTL show a worsening wind speed bias resulting from the suppression of extremes (Fig. 3e); however, there is no statistically significant difference between their predictions. This suggests that the degradation observed in GraphCast at full resolution, relative to to GDPS-CTL (Fig. 3a), is primarily due to fine-scale smoothing.
After filtering, regional deteriorations in GraphCast predictions are accentuated and regional improvements are significantly reduced, with the average reduction factor for $f_{RMSE}$ from full- to low-resolution verification being roughly between 3 to 4% lower than the values obtained at full resolution (c.f. the columns in Fig. 5). This further confirms that the lower effective resolution of GraphCast provides a significant advantage over the GDPS-CTL. Nevertheless, over most regions, and for both seasonal periods, the large scales in GraphCast forecasts are still, on average, considerably closer to tropospheric radiosonde observations than those from GDPS-CTL, implying that they contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
@@ 4.2 Implementation in GEM
To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales. For global simulations, such a filter should ideally employ a spherical harmonics-based spectral decomposition. However, this would require multiple transformations of model solutions between the model’s overlapping limited-area Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost. As the present study is primarily of a proof-of-concept nature, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids. The DCT-based spectral filter employs two cut-off wavelengths for large and small scales, denoted by $\lambda_{LS}$ and $\lambda_{SS}$, respectively. Scales larger than $\lambda_{LS}$ are fully retained by the filter, whereas scales smaller than $\lambda_{SS}$ are entirely removed, with a partial filtering of the scales in between (appendix C). This partial filtering, often referred to as soft cut-off (Husain et al., 2014), helps to minimize Gibbs oscillations (Sardeshmukh and Hoskins, 1984). The values of $\lambda_{LS}$ and $\lambda_{SS}$ are selected to produce a qualitatively acceptable response of the filter when tested over a global Gaussian grid of equivalent resolution (appendix C). Overall, the evaluation of $\left[F_{GC}-F_{GEM}\right]_{LS}$ for the purpose of spectral nudging involves computing the spectral coefficients of $(F_{GC}-F_{GEM})$ using a DCT, followed by the application of the spectral filter described in appendix C. An inverse DCT is then applied to obtain the large-scale differences in physical space.
@@ 4.3 Optimal configuration
A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for the hybrid GDPS forecasts, referred to as GDPS-SN. Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN, and GraphCast are all initialized with the same analyses.
@@ 4.3.3 Nudging length scales
It was shown earlier in section 3 that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs. 1 and 2). The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging. However, improvements with such an approach over stationary cut-off scales are found to be negligible. Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km. This implies that scales larger than 2750 km are entirely retained, while those smaller than 2250 km are completely eliminated. The scales in between are subjected to partial filtering (see Fig. 20). Over a global Gaussian grid of comparable resolution, the filter response is approximately equivalent to a soft cut-off between 5500-km and 2000-km scales (i.e., $\lambda_{LS}$=5500 km and $\lambda_{SS}$=2000 km), which provides a more relevant context for interpreting the results. This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs. 1 and 2).
@@ 5.1 Verification with global power spectra
The power spectra of the GDPS-SN forecast variables are indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can resolve the full range of scales present in the GDPS-CTL without significant smoothing. The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq. 1), for GDPS-SN also has values very close to the desired value of 1 for all prognostic variables, lead times, and spatial scales (Figs. 7a, b). The only exception is a 10% reduction of $\gamma$ for scales between 2000 km and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs. 7a, b). This reduction is explained by GraphCast having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4; Figs. 2a, b). Sensitivity tests with time-varying cut-off length scales ($\lambda_{LS}$ and $\lambda_{SS}$) that more strictly follow GraphCast’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL. This is because larger cut-offs imply a weaker GraphCast-based constraint at synoptic scales in the hybrid system.
@@
Comparison of the spectral coherence (Eq. 2) between the forecasts and GDPS analysis shows improved skill in GDPS-SN over GDPS-CTL (Figs. 7c, d). The improvements are somewhat smaller than those seen with GraphCast at shorter lead times (up to day 3), but become comparable thereafter (Figs 7c, d). The scales relevant for short-range improvements are smaller than those targeted by the chosen spectral nudging configuration. Inclusion of finer scales for nudging would improve short-range coherence, but it would result in problematic variance deficiency at longer leads.
Figure 7: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL. Additionally, for convenience of comparison, GraphCast has been added in Figs. 7c and d. The small- and large-scale cut-offs for the DCT-based filter over a global Gaussian grid (appendix C) are approximately 2000 km and 5500 km, respectively, as indicated with the shaded area.
@@
The impact of spectral nudging in the physical space is shown using maps of transient-eddy anomalies of temperature and specific humidity in Fig. 8. Nudging ensures that the large-scale temperature anomalies closely resemble GraphCast (Figs. 8a–c). Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are found to be more aligned with GraphCast than with GDPS-CTL (Figs. 8d–f). This reinforces the findings of previous studies on spectral nudging (Husain et al., 2014). Regarding fine scales, Fig. 8 shows that GDPS-SN has similar level of variability as GDPS-CTL in contrast to the heavily smoothed GraphCast fields.
@@
Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature (°$\degree$°C), and (d–f) specific humidity (g kg-1), valid at $0$ UTC, 18 February 2022 for GDPS-CTL (left), GraphCast (middle), and GDPS-SN (right). The black rectangles in Figs. 8 d-f indicate the region that is later used for presenting results in Fig. 9.
An illustrative example of the value added by the proposed hybrid system is presented in Fig. 9, which shows the 5.5-day forecasts by GDPS-CTL, GraphCast, and GDPS-SN, as well as GDPS analysis, valid at $0$ UTC on 18 February 2022. GDPS analysis shows the presence of the extratropical winter storm Eunice, which is also predicted by GDPS-SN, whereas both GDPS-CTL and GraphCast fail to predict the storm for this lead time. It is important to note that all three models are capable of predicting the storm at the next initialization time ($0$ UTC on 14 February 2022; not shown). These results imply that, for a lead time of 5.5 days, although GraphCast may predict large scales that are favorable for the formation of Eunice, it fails to develop the storm, presumably due to excessive smoothing at scales smaller than 2000 km. Since GDPS-SN does not suffer from similar smoothing, it can effectively leverage GraphCast’s more accurate large-scale information, resulting in a significant gain in predictability for this specific event, compared to both GraphCast and GDPS-CTL. More importantly, these results demonstrate that large-scale nudging of GDPS-SN toward GraphCast can lead to mesoscale features that are substantially different from both GDPS-CTL and GraphCast. In other words, the spectral nudging-based hybrid system has the potential to add substantial value beyond what is attainable through any offline post-processing method that combines NWP and AI models.
Figure 9: Maps of mean sea-level pressure (hPa) and 10-m wind speed (m s-1), valid at $0$ UTC, 18 February 2022 from: (a) GDPS analyses, and 5.5-day forecasts from (b) GDPS-CTL, (c) GraphCast, and (d) GDPS-SN. Results are presented over a domain identified with a black rectangle in Fig. 8d–f.